OpenStudy (anonymous):

z varies directly with x and inversely with y^2. When x=2 and y=5, z=8. What is the value of z when x=4 and y=9

OpenStudy (lgbasallote):

z varies directly with x and inversely with y^2 therefore \[z = \frac{kx}{y^2}\] when x = 2 and y = 5 then z = 8 \[8 = \frac{2k}{5^2}\] \[8 = \frac{2k}{25}\] \[200 = 2k\] \[k = 100\] when x = 4 and y = 9 \[z = \frac{kx}{y^2}\] \[z = \frac{(100)(4)}{9^2}\] \[z = \frac{400}{81}\]

OpenStudy (lgbasallote):

got that?

OpenStudy (anonymous):

Yea that makes sense! Thank you!

OpenStudy (lgbasallote):

<tips hat>